# Random covering by rectangles on self-similar carpets

### v1

@@ Abstract.
In thus article, given a base-$b$ self-similar set $K$, we study the random covering of $K$ by horizontal or vertical rectangles, with respect to the Alfhors-regular measure on $K$, and the rectangular shrinking target problem on $K$.
@@ 1. Introduction
where $i.o.$ (`` infinitely often ") means that the inequality holds for infinitely many $n$. Such problem were originally born in Diophantine approximation, where one studies sets of real numbers or vectors approximable by rational numbers or vectors at a given speed rate. Such questions are also natural in multifractal analysis, as for many mappings, the regularity at a given point depends of its rate of approximation by specific dyadic numbers or rational numbers (see [19] for instance) and in dynamical and random approximation, as, for instance, given an ergodic system $(T,\mu),$ the local dimension at a given point $x$ depends on its approximation rate by typical $\mu$ orbits, provided that $\mu$-mixes sufficiently fast (see [16, 10]). In particular, the theory of random approximation as raised many interests these last 30 years and has known recently many developments. Given $\mu\in\mathcal{M}(\mathbb{R}^{d})$ a probability measure and $(X_{n})_{n\in\mathbb{N}}$ an i.i.d. sequence of random variable of law $\mu$ it was established in [20] that, denoting $\overline{\dim}_{H}\mu$ the upper Hausdorff dimension of $\mu$ (see Definition 2.2), for any $\delta\geq\frac{1}{\overline{\dim}_{H}\mu},$ almost surely, one has
@@
where $\mathcal{H}^{t}_{\infty}(O_{n})$ denotes the Hausdorff content of dimension $s$ of $O_{n}$ (Definition 5).
@@
where $\pi_{2}\mu_{0}$ denotes the projection of $\mu_{0}$ along the $y$-axis and $D_{\pi_{2}\mu_{0}}$ its multifractal spectrum (Definition 2.4). Let $\beta_{\tau_{1},\tau_{2}}$ be the smallest solution (when well defined) of $v_{\tau}(\beta)=\frac{1}{\tau}.$
@@
For a more precise statement, we refer to Theorem 3.1 below.
@@
The shrinking target problem on fractals has known many developments since, one refers for instance to [3, 1, 7, 4] In the present article, we study the shrinking target when ``the targets " are taken to be rectangles rather than balls. Unlike the case of balls, our result will depend, in general, on the choice of $x$, the center of our targets. More precisely, let $\nu$ be a $\times b$-ergodic measure supported on $K$, let $1\leq\tau_{1}\leq\tau_{2}$ be two real numbers and write
$V_{\tau_{1},\tau_{2}}=\limsup_{n\geq 1,(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}\Big(f_{i_{1}}\circ...\circ f_{i_{n}}(x)+(-b^{-\tau_{1}n},b^{-\tau_{1}n})\times(-b^{-\tau_{2}n},b^{-\tau_{2}n})\Big).$
@@
$V_{\tau_{1},\tau_{2}}=\left\{y\in K:\ T^{n}(y)\in\Big(x+(-b^{-(\tau_{1}-1)n},b^{-(\tau_{1}-1)n})\times(-b^{-(\tau_{2}-1)n},b^{-(\tau_{2}-1)n})\Big)\right\}.$
@@
where $\alpha_{\nu}$ is the almost sure local dimension of $\pi_{2}(x)$ with respect to $\pi_{2}\mu_{0}$ (see Proposition 3.3). We refer to Theorem 3.4 for a more general statement and Corollary 3.5 for a formula for general approximation function along the $x$ and $y$-axis. Notice that set of possible dimensions (which are all attained) for $V_{\tau_{1},\tau_{2}}$ when $\nu$ varies in the set of ergodic measures is
@@ 2.3. Multifractal analysis of self-similar measure satisfying OSC
  • the mapping $\alpha\mapsto D_{\mu}(\alpha)$ satisfies is concave and reaches its maximum on $\alpha=\dim_{H}\mu,$
@@
In addition of these properties, when $\mu$ is a self-similar measure satisfying the open set condition, it is also known that the multifractal spectrum and the so-called coarce multifractal spectrum coincides. More precisely, we have the following large deviation estimates.
@@
where $\mathcal{T}$ is a $r$-packing of $\mathrm{supp}(\mu)$ with, for every $B\in\mathcal{T}$,
@@ 3.2. Tree approximation
Consider again $S=\left\{f_{1},..,f_{m}\right\}$, a two dimensional base-$b$ missing digit IFS of attractor $K$ and $\mu_{0}$ the Alfhors-regular self-similar measure on $K$. The following proposition is necessary to state our main result and will be established in the next section, as Proposition 4.1 applied with $\pi_{2}\nu$ (which can be identified with a $\times-b$ ergodic measure on $\mathbb{T}^{1}$).
@@
Then, for every $1\leq\tau_{1}\leq\tau_{2},$ for any $\times b$-ergodic measure $\nu$ with $\mathrm{supp}(\nu)\subset K$, writing
@@
$\dim_{H}\limsup_{\underline{i}\in\bigcup_{k\geq 1}\left\{1,...,m\right\}^{k}}\Big(f_{\underline{i}}(x)+(-\psi(|\underline{i}|),\psi(|\underline{i}|)\times(-\theta(|\underline{i}|),\theta(|\underline{i}|))\Big)=\max\left\{g_{1}(\psi,\theta,\nu),g_{2}(\psi,\theta,\nu)\right\}.$
@@
As mentioned in introduction, Corollary 3.5 was also obtained by Allen, Jordan and Ward in [2], using a different method, under the weaker assumption that the coding of $x=(x_{n}:=(x_{n}^{1},x_{n}^{2}))_{n\in\mathbb{N}}\in\mathcal{A}^{\mathbb{N}}$ satisfies that, writing $S=S_{\mathcal{A}},$ for each $0\leq j\leq b-1$ for which there exists $0\leq i\leq b-1$ so that $(i,j)\in\mathcal{A},$ one has
@@
The next three sections are dedicated to the proofs of Theorem 3.1 and Theorem 3.4. The Section 4 establishes some useful general results regarding $\times-b$ ergodic measures, provides some recall on the mass transference principle for self-similar measures and establishes important content estimates, useful to the proof of both theorems. In the last section, we finally prove Theorem 3.1 and Theorem 3.4.
@@
4.1. Results regarding $\times-b$ ergodic measures
Let us fix $b\in\mathbb{N}$ and $\eta$ a $\times b$-ergodic measure on $\mathbb{T}$. For $0\leq i\leq b-1,$ let us denote $f_{i}$ the mapping defined for every $x\in\mathbb{T}$ by $f_{i}(x)=\frac{x+i}{b}$ and $S_{b}=\left\{f_{i}\right\}_{0\leq i\leq b-1}.$ The goal of this section is to prove the following result.
@@ 4.1. Results regarding $\times-b$ ergodic measures
4.2. Recall on mass transference principle for self-similar measures
@@ 4.3. Content estimates
Let $\nu\in\mathcal{M}([0,1]^{2})$ be a $\times$-b ergodic measure. Then, for $\nu$-almost every $x$ for every $\varepsilon>0$, there exists $r_{x}>0$ such that for every $r\leq r_{x}$, for every $0\leq s\leq s_{0},$ one has
@@
Fix $\varepsilon_{0}>0.$ Recall that, since $\pi_{2}\nu$ is $\times-b$ ergodic, either $\nu$ is a periodic measure (hence is carried by finitely many atoms) or $\pi_{2}\nu$ is diffuse, which case, by Corollary 4.4, there exists $r_{1}>0$ small enough so that for every $r\leq r_{1},$ one has
@@
Case 3: $\frac{1}{\tau}\geq v_{\tau}(\kappa_{2}).$

### v2

@@ Abstract.
In this article, given a base-$b$ self-similar set $K$, we study the random covering of $K$ by horizontal or vertical rectangles, with respect to the Alfhors-regular measure on $K$, and the rectangular shrinking target problem on $K$.
@@ 1. Introduction
where $i.o.$ (`` infinitely often ") means that the inequality holds for infinitely many $n$. Such problems were originally born in Diophantine approximation, where one studies sets of real numbers or vectors approximable by rational numbers or vectors at a given speed rate. These questions are also natural in multifractal analysis, as for many mappings, the regularity at a given point depends on its rate of approximation by specific dyadic numbers or rational numbers (see [19] for instance) and in dynamical and random approximation, as, for instance, given an ergodic system $(T,\mu),$ the local dimension at a given point $x$ depends on its approximation rate by typical $\mu$ orbits, provided that $\mu$ mixes sufficiently fast (see [16, 10]). In particular the theory of random approximation as raised many interests these last 30 years and has known recently many developments. Given $\mu\in\mathcal{M}(\mathbb{R}^{d})$ a probability measure and $(X_{n})_{n\in\mathbb{N}}$ an i.i.d. sequence of random variable of law $\mu$ it was established in [20] that, denoting $\overline{\dim}_{H}\mu$ the upper Hausdorff dimension of $\mu$ (see Definition 2.2), for any $\delta\geq\frac{1}{\overline{\dim}_{H}\mu},$ almost surely, one has
@@
where $\mathcal{H}^{t}_{\infty}(O_{n})$ denotes the Hausdorff content of dimension $t$ of $O_{n}$ (Definition 5).
@@
where $\pi_{2}\mu_{0}$ denotes the projection of $\mu_{0}$ along the $y$-axis and $D_{\pi_{2}\mu_{0}}$ its multifractal spectrum (Definition 2.4). Let $\beta_{\tau_{1},\tau_{2}}$ be the smallest solution (when well defined) of $v_{\tau}(\beta)=\frac{1}{\tau_{1}}.$
@@
For a more precise statement (in particular regarding the value of $\kappa_{2}$), we refer to Theorem 3.1 below.
@@
The shrinking target problem on fractals has known many developments since and an interested reader may refer to [3, 1, 7, 4] for various related results. In the present article, we study the shrinking target when ``the targets " are taken to be rectangles rather than balls. Unlike the case of balls, our result will depend, in general, on the choice of $x\in K$, the center of our targets. More precisely, let $\nu$ be a $\times b$-ergodic measure supported on $K$, let $1\leq\tau_{1}\leq\tau_{2}$ be two real numbers and write
$V_{\tau_{1},\tau_{2}}(x)=\limsup_{n\geq 1,(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}\Big(f_{i_{1}}\circ...\circ f_{i_{n}}(x)+(-b^{-\tau_{1}n},b^{-\tau_{1}n})\times(-b^{-\tau_{2}n},b^{-\tau_{2}n})\Big).$
@@
$V_{\tau_{1},\tau_{2}}(x)=\left\{y\in K:\ T^{n}(y)\in\Big(x+(-b^{-(\tau_{1}-1)n},b^{-(\tau_{1}-1)n})\times(-b^{-(\tau_{2}-1)n},b^{-(\tau_{2}-1)n})\Big)\right\}.$
@@
where $\alpha_{\nu}$ is the almost sure local dimension of $\pi_{2}(x)$ with respect to $\pi_{2}\mu_{0}$ (see Proposition 3.3). We refer to Theorem 3.4 for a more general statement and Corollary 3.5 for a formula holding for general approximation function along the $x$ and $y$-axis. Notice that set of possible dimensions (which are all attained) for $V_{\tau_{1},\tau_{2}}$ when $\nu$ varies in the set of ergodic measures is
@@ 2.3. Multifractal analysis of self-similar measure satisfying OSC
  • the mapping $\alpha\mapsto D_{\mu}(\alpha)$ is concave and reaches its maximum on $\alpha\geq\dim_{H}\mu$ such that $D_{\mu}(\alpha)=\dim_{H}K.$
@@
In addition of these properties, when $\mu$ is a self-similar measure satisfying the open set condition, it is also known that the multifractal spectrum and the so-called coarse multifractal spectrum coincides. More precisely, we have the following large deviation estimates.
@@
where $\mathcal{T}$ is a maximal $r$-packing of $\mathrm{supp}(\mu)$ with, for every $B\in\mathcal{T}$,
@@ 3.2. Tree approximation
Consider again $S=\left\{f_{1},..,f_{m}\right\}$, a two dimensional base-$b$ missing digit IFS of attractor $K$ and $\mu_{0}$ the Alfhors-regular self-similar measure on $K$. The following proposition is necessary to state our main result and will be established in the next section, as Proposition 4.1 applied with $\pi_{2}\nu$ (which can be identified with a $\times b$ ergodic measure on $\mathbb{T}^{1}$).
@@
In the next theorem, given a word $\underline{i}=(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n},$ one writes
$f_{\underline{i}}=f_{i_{1}}\circ...\circ f_{i_{n}}.$
@@
Then, for every $1\leq\tau_{1}\leq\tau_{2},$ for any $\times b$ ergodic measure $\nu$ with $\mathrm{supp}(\nu)\subset K$, writing
@@
$\dim_{H}\limsup_{\underline{i}\in\bigcup_{k\geq 1}\left\{1,...,m\right\}^{k}}\Big(f_{\underline{i}}(x)+(-\psi(|\underline{i}|),\psi(|\underline{i}|)\times(-\theta(|\underline{i}|),\theta(|\underline{i}|))\Big)$
$=\max\left\{g_{1}(\psi,\theta,\nu),g_{2}(\psi,\theta,\nu)\right\}.$
@@
As mentioned in the introduction, Corollary 3.5 was also obtained by Allen, Jordan and Ward in [2], using a different method, under the weaker assumption that the coding of $x=(x_{n}:=(x_{n}^{1},x_{n}^{2}))_{n\in\mathbb{N}}\in\mathcal{A}^{\mathbb{N}}$ satisfies that, writing $S=S_{\mathcal{A}},$ for each $0\leq j\leq b-1$ for which there exists $0\leq i\leq b-1$ so that $(i,j)\in\mathcal{A},$ one has
@@
The next three sections are dedicated to the proofs of Theorem 3.1 and Theorem 3.4. The Section 4 establishes some useful general results regarding $\times b$ ergodic measures, provides some recalls on the mass transference principle for self-similar measures and establishes important content estimates, useful to the proof of both theorems. In the last section, we finally prove Theorem 3.1 and Theorem 3.4.
@@
4.1. Results regarding $\times b$ ergodic measures
Let us fix $b\in\mathbb{N}$ and $\eta$ a $\times b$ ergodic measure on $\mathbb{T}$. For $0\leq i\leq b-1,$ let us denote $f_{i}$ the mapping defined for every $x\in\mathbb{T}$ by $f_{i}(x)=\frac{x+i}{b}$ and $S_{b}=\left\{f_{i}\right\}_{0\leq i\leq b-1}.$ The goal of this section is to prove the following result.
@@ 4.1. Results regarding $\times b$ ergodic measures
4.2. Recalls on mass transference principle for self-similar measures
@@ 4.3. Content estimates
Let $\nu\in\mathcal{M}([0,1]^{2})$ be a $\times b$ ergodic measure. Then, for $\nu$-almost every $x$ for every $\varepsilon>0$, there exists $r_{x}>0$ such that for every $r\leq r_{x}$, for every $0\leq s\leq s_{0},$ one has
@@
Fix $\varepsilon_{0}>0.$ Recall that, since $\pi_{2}\nu$ is $\times b$ ergodic, either $\nu$ is a periodic measure (hence is carried by finitely many atoms) or $\pi_{2}\nu$ is diffuse, which case, by Corollary 4.4, there exists $r_{1}>0$ small enough so that for every $r\leq r_{1},$ one has
@@
Case 3: $\frac{1}{\tau_{1}}\geq v_{\tau}(\kappa_{2}).$